A man has five sons, each three years older than the next younger. The age of the oldest in three years will be three times the present age of the second youngest. How old is each son from youngest to oldest?
(Not from a text) seems simple. Would the youngest be 3 and the oldest 15?
Thank you for your help.

Found 2 solutions by checkley71, Paul:
Answer by checkley71(8403)

YES. (X+3)*3=(X+12)+3 OR 3X+9=X+15 OR 2X=6 OR X=3 THUS THE YOPUNGEST IS 3 YEARS OLD THE SECOND YOUNGEST IS 6 AND THE OLDEST IS 15 THUS WE HAVE 3,6,9,12&15 & IN 3 YEARS THE OLDEST WILL BE 18 WHICH IS 3*6=18.

Answer by Paul(988)

5 sons --> x, x+3, x+6, x+9, x+12
Age of the oldest in 3 years = x+12+3=x+15
Equation:
x+15=3(x+3)
x+15=3x+9
-2x=-6
x=3
Hence, the yougest son is 3 (add 3 to each of the other sons ages)
Paul.